OpenStudy (anonymous):

Hi, can someone explain and demonstrate how to solve this problem please? 578 = (346.7/346.7-x)551

OpenStudy (anonymous):

\[ 578 = \frac{346.7}{346.7-x}551\]

OpenStudy (anonymous):

not sure where the 551 is, numerator or denominator

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

551 was just multiplied by the outside of the fraction, which was in parentheses. Does that help at all?

OpenStudy (amistre64):

551 like that tends to be exponent

OpenStudy (amistre64):

so are we like coming in at the tail end of a problem that was half solved and we have to determine what went wrong?

OpenStudy (anonymous):

no, that's just the full problem. 551 has a multiplication symbol before it, so i don't believe it's an exponent. :/

OpenStudy (amistre64):

ahhh, ok. we get typos in here alot that suggest that this configuration ...... like x2 is somehow x^2 :)

OpenStudy (anonymous):

578(346.7-x)=346.7(551) cross multiplying 200392.6-578x=191031.7 multiplying 9360.9=578x x=16.19(ans)

OpenStudy (oaktree):

Yup, I got that too.

OpenStudy (anonymous):

ok guys thanks!! I believe this is the correct answer as well. I just wasn't sure how to get that subtracted x out of the fraction. thanks for all your help!

OpenStudy (anonymous):

no mention :)